# Wigner function method for the Gibbons–Hawking and the Unruh effect

### v1

@@ 1 Introduction
Imagine an observer at rest with the empty space of the expanding universe (Fig. 1). Suppose the observer is equipped with a sensitive spectrometer for detecting electromagnetic radiation. The spectrometer picks up the noise of vacuum fluctuations and also some extra noise that closely resembles the Hawking radiation of black holes [1, 2]. The extra noise depends on the expansion, and it is due to a difference in the measure of time the observer and the vacuum perceive. The spectrometer decomposes the field into frequency components with respect to the proper time of the observer, whereas the fluctuations of the field (Fig. 1) are continuously red–shifted by cosmic expansion [3]. A similar difference in time appears in the opposite scenario: an accelerated observer in a vacuum at rest [4, 5, 6] (in non–expanding Minkowski space). There the proper time of the observer differs from the time the vacuum experiences, too. The extra radiation is related to the Hawking radiation of black holes [1], because in both cases the observer is shrouded by horizons. In the expanding universe, it is the cosmological horizon [3] where the expansion velocity reaches the speed of light. In the case of the accelerated observer, it is the Rindler horizon [1] where the observer approaches the speed of light. For an exponentially expanding, spatially flat universe (de–Sitter space) the cosmological horizon is light–like [2] as the event horizon of an ideal black hole [1] and the observer would measure a perfect Planck spectrum [2]. But the real evolution of the universe is not exponential yet [7]. What is the spectrum there? The same question applies to the accelerated observer. A uniformly accelerated observer would measure a true Planck spectrum. But real accelerations must begin and end some time or may not be uniform at all. Is then the spectrum still Planckian?
@@
In this paper, we hope to shed some light on the radiation spectrum in the expanding universe, known as the Gibbons–Hawking effect [2], for non–exponential expansion and on the radiation perceived by accelerated observers, the Fulling–Davies–Unruh effect [4, 5, 6] or Unruh effect in short, for non–uniform acceleration. The paper builds upon and corrects previous work [23, 26, 27] using the Wigner function [28, 29, 30] for defining and calculating time–dependent spectra. We found excellent agreement with a Planck spectrum for one class of models, and poor agreement for another. What the condition for thermality and the effective temperatures are remains a mystery; it would probably take someone like Renaud Parentani to clarify it.
@@ 2 Correlation function
Consider a spatially flat universe expanding with scale factor $a(t)$ where $t$ denotes the time perceived by an observer at rest, the cosmological time. The quantum vacuum propagates with conformal time [7]:
@@
The difference between the time $t$ of the observer and the time $\tau$ of the vacuum is the key to the Gibbons–Hawking effect [27]. For simplicity, we treat the electromagnetic field as a conformally coupled scalar field [27] $\widehat{A}$ with the correlation function
$K=\frac{1}{2}\langle 0|\widehat{A}_{1}\widehat{A}_{2}+\widehat{A}_{2}\widehat{A}_{1}|0\rangle$ (2)
of vacuum noise at the two space–time points $\{t_{1},{\bm{r}}_{1}\}$ and $\{t_{2},{\bm{r}}_{2}\}$ in co–moving coordinates indicated by the indices in $\widehat{A}_{1}$ and $\widehat{A}_{2}$. The vacuum energy density and pressure of the field can be calculated from $K$ by differentiating with respect to the space–time coordinates. Note that such a correlation function has been measured inside a dielectric material [32]. The correlation $K$ is the real part of the analytic function $\langle 0|\widehat{A}_{1}\widehat{A}_{2}|0\rangle$ while its imaginary part corresponds to the dissipation [31], the difference between the retarded and the advanced Green functions $G_{\pm}=-(4\pi r)^{-1}\delta(\tau_{21}\mp r/c)$ with $\tau_{21}=\tau_{2}-\tau_{1}$ and $r=|{\bm{r}}_{2}-{\bm{r}}_{1}|$. The fluctuation–dissipation theorem [31] relates the two, turning each delta function in the dissipation to a pole in the correlation. One obtains [23]:
$K=-\frac{1}{(2\pi)^{2}s^{2}}$ (3)
with
@@
with constant coefficients $\alpha$, $\beta$, $\gamma$, $\delta$ the correlation function (5) remains unchanged.
@@
One gets from the fluctuation–dissipation theorem with Kubo–Martin–Schwinger relation [31] the correlation function:
@@
Finally, we turn from the Gibbons–Hawking effect of the observer at rest with the expanding universe to the Unruh effect of the accelerated observer. In the Unruh effect [4, 5, 6] a detector moves in empty Minkowski space on the space–time trajectory $t(\tau)$ and ${\bm{r}}(\tau)$ parameterized by the proper time of the detector, $\tau=\int\mathrm{d}\tau$ with $\mathrm{d}\tau^{2}=\mathrm{d}t^{2}-\mathrm{d}{\bm{r}}^{2}/c^{2}$. Here the roles of $t$ and $\tau$ are reversed: the vacuum propagates with Minkowski time $t$ while the spectrometer of the detector responds to the proper time $\tau$. The detector thus samples the correlation function (3) with
@@ 3 Wigner function
and the same expression with $\omega$ replaced by $|\omega|$ for $\omega<0$. We may interpret this Wigner function as being proportional to the thermal energy of a bosonic point particle: the $\omega/2$ term describes the zero–point energy and the other term the Planck distribution of a thermal state with temperature (8). The point detector acts as a point particle in thermal equilibrium with Gibbons–Hawking temperature (8). Mathematically, the zero–point term appears due to the double pole of the correlation function (5) at $\theta=0$. We may remove this term by moving the integration contour in formula (11) infinitesimally above the real axis (Fig. 2) for $\omega>0$ (or below it for $\omega<0$).
This procedure has another advantage [27]. At the beginning of time, say at $t=0$, the scale factor develops a branch point [41] for a realistic universe (such that $a$ is no longer real or positive for $t<0$ — there was no discernible universe before the beginning of time. Suppose that $a\propto t^{1/\gamma}$ with constant $\gamma\geq 1$ for $t\sim 0$. From this follows [Eq. (5)] that $K\propto(t+\theta/2)^{-1/\gamma}+(t-\theta/2)^{-1/\gamma}$ as leading contribution near $\theta=\pm 2t$. We get from Wigner’s formula (11):
@@
Fourier transformations generate oscillations near sharp features; for the expanding universe we get spectral oscillations from the branch point of $a(t)$ at the beginning of time. Moving the integration contour slightly above or below the real axis avoids the branch points (Fig. 2) and thus removes this artefact. In a previous study [23] the integration was left on the real axis, which would have produced highly oscillatory, meaningless spectra if they were calculated.
@@
The constant $H_{\infty}$ amounts to the Hubble parameter infinitely in the future when $a\rightarrow\infty$. For $w=0$ and hence $\gamma=3/2$ this model describes the current era of the universe turning from matter to vacuum domination. We obtain from the definition (16) of the Hubble parameter:
@@
For $H_{0}t\ll-1$ the conformal time approaches $t/a_{0}$ and for $H_{0}t\gg+1$ it tends to $t/a_{1}$, both as it should for constant scale factors $a$ according to definition (1). Figure 5b shows the Wigner function. For $H_{0}t\lesssim-1$ we get a good Planck spectrum, but for $H_{0}t\gtrsim+1$ the Wigner function becomes negative. In quantum optics [30], negative regions of the Wigner function indicate nonclassical features [40]; we are not sure what they mean here.
@@ 4 Conclusions
We applied an amended [27] Wigner function [28, 29, 30] for calculating the spectrum of vacuum fluctuations for an observer at rest in the expanding universe or, equivalently, for an accelerated observer in a space at rest (in empty Minkowski space). The literature mostly focuses on exponential expansion (de–Sitter space) and uniform acceleration (Rindler trajectories) because these are the ideal cases where the cosmological or the Rindler horizons are genuine event horizons. Here we have developed tools for addressing the real cases of non–exponential expansion and non–uniform acceleration. The key feature that allows the calculation of physically meaningful spectra is the modification of the integration contour in the Wigner function [27]. Otherwise the spectra would be highly oscillatory and meaningless. For a class of models that includes the model describing the current cosmological era, we found excellent fits with Planck spectra. The vacuum noise in the expanding universe appears as thermal noise. However, we were not able to derive a simple analytic formula for the radiation temperature; a previous formula [27] for the high–frequency limit turned out to be incorrect. For high frequencies in the exponential tail of the spectrum, the Wigner function appears to exhibit spurious oscillations, and the Wigner method seems to fail there. For another type of cosmic evolution that can be experimentally tested in laboratory analogues [14, 15, 16, 17] we found non–Planckian, negative Wigner functions. In quantum optics [30] negative Wigner functions indicate non–classical features. What are they in our case? When is the spectrum Planckian? What is the temperature? When science answers one question, many more questions arise. We leave it to future readers to find some of the answers.

### v2

@@ 1 Introduction
Imagine an observer at rest with the empty space of the expanding universe (Fig. 1). Suppose the observer is equipped with a sensitive spectrometer for detecting electromagnetic radiation. The spectrometer picks up the noise of vacuum fluctuations and also some extra noise that closely resembles the Hawking radiation of black holes [1, 2]. The extra noise depends on the expansion, and it is due to a difference in the measure of time the observer and the vacuum perceive. The spectrometer decomposes the field into frequency components with respect to the proper time of the observer, whereas the fluctuations of the field (Fig. 1) are continuously red–shifted by cosmic expansion [3]. A similar difference in time appears in the opposite scenario: an accelerated observer in a vacuum at rest [4, 5, 6, 7] (in non–expanding Minkowski space). There the proper time of the observer differs from the time the vacuum experiences, too. The extra radiation is related to the Hawking radiation of black holes [1], because in both cases the observer is shrouded by horizons. In the expanding universe, it is the cosmological horizon [3] where the expansion velocity reaches the speed of light. In the case of the accelerated observer, it is the Rindler horizon [1] where the observer approaches the speed of light. For an exponentially expanding, spatially flat universe (de–Sitter space) the cosmological horizon is light–like [2] as the event horizon of an ideal black hole [1] and the observer would measure a perfect Planck spectrum [2]. But the real evolution of the universe is not exponential yet [8]. What is the spectrum there? The same question applies to the accelerated observer. A uniformly accelerated observer would measure a true Planck spectrum. But real accelerations must begin and end some time or may not be uniform at all. Is then the spectrum still Planckian?
@@
In this paper, we hope to shed some light on the radiation spectrum in the expanding universe, known as the Gibbons–Hawking effect [2], for non–exponential expansion and on the radiation perceived by accelerated observers, the Fulling–Davies–Unruh effect [4, 5, 6, 7] or Unruh effect in short, for non–uniform acceleration. The paper builds upon and corrects previous work [24, 27, 28] using the Wigner function [31, 32, 33] for defining and calculating time–dependent spectra. We found excellent agreement with a Planck spectrum for one class of models, and poor agreement for another. What the condition for thermality and the effective temperatures are remains a mystery; it would probably take someone like Renaud Parentani to clarify it.
@@ 2 Correlation function
Consider a spatially flat universe expanding with scale factor $a(t)$ where $t$ denotes the time perceived by an observer at rest, the cosmological time. The quantum vacuum of the electromagnetic field propagates with conformal time [8] (Appendix A):
@@
The difference between the time $t$ of the observer and the time $\tau$ of the vacuum is the key to the Gibbons–Hawking effect [28].Consider the correlation function of the electromagnetic vector potential at the two space–time points $\{t_{1},{\bm{r}}_{1}\}$ and $\{t_{2},{\bm{r}}_{2}\}$ in co–moving coordinates [8].In an isotropic and homogeneous universe [8], the two electromagnetic polarizations are independent from each other [34] and have equal correlation functions:
$K_{0}=\frac{\varepsilon_{0}c}{2\hbar}\langle 0|\widehat{A}_{1}\widehat{A}_{2}+\widehat{A}_{2}\widehat{A}_{1}|0\rangle$ (2)
where $\varepsilon_{0}$ denotes the electric permittivity of the vacuum ($c$ the speed of light and $\hbar$ the reduced Planck constant) while the indices in $\widehat{A}_{1}$ and $\widehat{A}_{2}$ refer to the space–time points $\{t_{1},{\bm{r}}_{1}\}$ and $\{t_{2},{\bm{r}}_{2}\}$ the field operators are evaluated at. The vacuum energy density and pressure of the field can be calculated from $K_{0}$ by differentiating with respect to the space–time coordinates. The correlation $K_{0}$ is the real part of the analytic function $(\varepsilon_{0}c/\hbar)\langle 0|\widehat{A}_{1}\widehat{A}_{2}|0\rangle$ while its imaginary part corresponds to the dissipation [34]:
$\Gamma_{0}=\frac{\varepsilon_{0}c}{2\mathrm{i}\hbar}\,\big{[}\widehat{A}_{2},\widehat{A}_{1}\big{]}=\frac{1}{2c}(G_{+}-G_{-})$ (3)
where the $G_{\pm}$ are the retarded and advanced Green functions. These Green functions are entirely classical and independent of the quantum state. The advanced Green function $G_{-}$ describes a wave focusing at a point drain with unit strength, while the retarded Green function $G_{+}$ describes the wave emitted from a point source of unit strength. The difference between the advanced and the retarded Green function accounts for the dissipation. For analytic functions $f$ with
$K=\mathrm{Re}f\,,\quad\Gamma=\mathrm{Im}f$ (4)
the correlation $K$ of the fluctuations is related to the dissipation $\Gamma$. This is the fluctuation–dissipation theorem [34]. Note that the correlation $K_{0}$ depends on the quantum state, the vacuum state in our case, while the dissipation is state–independent. For the electromagnetic vacuum, we encode the state requiring that the Fourier transform of $f_{0}$ with respect to conformal time (1) contains only positive frequencies. This is the condition of the conformal vacuum and it comes from the fact that the modes of the electromagnetic field are waves of positive frequencies with respect to $\tau$. From this condition follows the analyticity and decay of $(\varepsilon_{0}c/\hbar)\langle 0|\widehat{A}_{1}\widehat{A}_{2}|0\rangle$ on the lower half $\tau$ plane and hence the Kramers–Kronig relation (known in mathematics as the Hilbert transformation):
$K_{0}(\tau)=-\frac{1}{\pi}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{-\infty}^{+\infty}\frac{\Gamma_{0}(\tau_{0})}{\tau_{0}-\tau}\,\mathrm{d}\tau_{0}\,.$ (5)
In a spatially flat universe, electromagnetic fields propagate with respect to conformal time (1) like in Minkowski space [8]. We thus have $G_{\pm}=-(4\pi r)^{-1}\delta(\tau_{21}\mp r/c)$ with $\tau_{21}=\tau_{2}-\tau_{1}$ and $r=|{\bm{r}}_{2}-{\bm{r}}_{1}|$. The Kramers–Kronig relation (5) turns each delta function in the dissipation to a pole in the correlation:
$K_{0}=-\frac{1}{8\pi^{2}r}\left(\frac{1}{c\tau_{21}-r}-\frac{1}{c\tau_{21}+r}\right).$ (6)
This is the correlation function of the electromagnetic vacuum fluctuations (Fig. 1). Note that such a correlation function has been measured inside a dielectric material [35]. If we divide the vector–potential component $A$ of each polarization by the scale factor $a$ we obtain the conformally–coupled scalar field $\phi$ (Appendix A). It is the correlations of this field we are considering here:
$K=\frac{K_{0}}{a_{1}a_{2}}=-\frac{1}{(2\pi)^{2}s^{2}}$ (7)
where we get from Eq. (6):
@@
with constant coefficients $\alpha$, $\beta$, $\gamma$, $\delta$ the correlation function (9) remains unchanged.Note that this feature is not related to Möbius transformations of the cosmological time $t$ studied elsewhere [29, 30] (despite using there the notation $\tau$ for $t$).
@@
One gets from the fluctuation–dissipation theorem (4) with Kubo–Martin–Schwinger relation [34] the correlation function:
@@
Finally, we turn from the Gibbons–Hawking effect of the observer at rest with the expanding universe to the Unruh effect of the accelerated observer. In the Unruh effect [4, 5, 6, 7] a detector moves in empty Minkowski space on the space–time trajectory $t(\tau)$ and ${\bm{r}}(\tau)$ parameterized by the proper time of the detector, $\tau=\int\mathrm{d}\tau$ with $\mathrm{d}\tau^{2}=\mathrm{d}t^{2}-\mathrm{d}{\bm{r}}^{2}/c^{2}$. Here the roles of $t$ and $\tau$ are reversed: the vacuum propagates with Minkowski time $t$ while the spectrometer of the detector responds to the proper time $\tau$. The detector thus samples the correlation function (7) with
@@ 3 Wigner function
independent of time $t$, and the same expression with $\omega$ replaced by $|\omega|$ for $\omega<0$. We may interpret this Wigner function as being proportional to the thermal energy of a bosonic point particle: the $\omega/2$ term describes the zero–point energy and the other term the Planck distribution of a thermal state with temperature (12). The point detector acts as a point particle in thermal equilibrium with Gibbons–Hawking temperature (12). Mathematically, the zero–point term appears due to the double pole of the correlation function (9) at $\theta=0$:applying for $r\neq 0$ Cauchy’s theorem through the single poles on the real $\theta$ axis and taking the limit $r\rightarrow 0$ produces this term. This is true not only for exponential expansion, but for any $a(t)$, which is consistent with the physical picture of this term describing the vacuum noise floor of the detector. We may remove the vacuum term by moving the integration contour in formula (15) infinitesimally above the real axis (Fig. 2) for $\omega>0$ (or below it for $\omega<0$) as this contour does not cross the poles and hence does not produce the vacuum term.
Moving the integration contour has another advantage [28]. At the beginning of time, say at $t=0$, the scale factor must develop a branch point [44] for a realistic universe such that $a$ is no longer real or positive for $t<0$ — there was no discernible universe before the beginning of time. This would restrict the integration interval in the Wigner function (15) to $[-2t,+2t]$. However, we may analytically continue the correlation function $K$ beyond the physically relevant interval $[-2t,+2t]$. As the spectrum should primarily depend on cosmologically short times around $t$ this extension should not affect it much. But there is a problem. Suppose that $a\propto t^{1/\gamma}$ with constant $\gamma\geq 1$ for $t\sim 0$. From this follows [Eq. (9)] that $K\propto(t+\theta/2)^{-1/\gamma}+(t-\theta/2)^{-1/\gamma}$ as leading contribution near $\theta=\pm 2t$. We obtain from Wigner’s formula (15):
@@
We thus get spectral oscillations. Fourier transformations are known to generate oscillations near sharp features; in our casethese originate from the branch points in the integrand of the Wigner function. Moving the integration contour slightly above or below the real axis avoids the branch points (Fig. 2) and thus removes this artefact. In a previous study [24] the integration was left on the real axis, which would have produced highly oscillatory, meaningless spectra if they were actually calculated, whereas for our integration contour we may get physically meaningful spectra of the extra noise due to cosmic expansion.
@@
The constant $H_{\infty}$ amounts to the Hubble parameter infinitely in the future when $a\rightarrow\infty$. For $w=0$ and hence $\gamma=3/2$ this model describes the current era of the universe turning from matter to vacuum domination. We obtain from Eq. (1) for the conformal time and definition (20) of the Hubble parameter:
@@
For $H_{0}t\ll-1$ the conformal time approaches $t/a_{0}$ and for $H_{0}t\gg+1$ it tends to $t/a_{1}$, both as it should for constant scale factors $a$, according to definition (1). Figure 5b shows the Wigner function. For $H_{0}t\lesssim-1$ we get a good Planck spectrum, but for $H_{0}t\gtrsim+1$ the Wigner function becomes negative. In quantum optics [33], negative regions of the Wigner function indicate nonclassical features [43]; we are not sure what they mean here.
@@ 4 Conclusions
We applied an amended [28] Wigner function [31, 32, 33] for calculating the spectrum of vacuum fluctuations for an observer at rest in the expanding universe or, equivalently, for an accelerated observer in a space at rest (in empty Minkowski space). The literature mostly focuses on exponential expansion (de–Sitter space) and uniform acceleration (Rindler trajectories) because these are the ideal cases where the cosmological or the Rindler horizons are genuine event horizons. Here we have developed tools for addressing the real cases of non–exponential expansion and non–uniform acceleration. The key feature that allows the calculation of physically meaningful spectra is the modification of the integration contour in the Wigner function [28]. Otherwise the spectra would be highly oscillatory and meaningless. For a class of models that includes the model describing the current cosmological era, we found excellent fits with Planck spectra. The vacuum noise in the expanding universe appears as thermal noise. However, we were not able to derive a simple analytic formula for the radiation temperature; a previous formula [28] for the high–frequency limit turned out to be incorrect. For high frequencies in the exponential tail of the spectrum, the Wigner function exhibits spurious oscillations, and the Wigner method seems to fail there. For another type of cosmic evolution that can be experimentally tested in laboratory analogues [15, 16, 17, 18] we found non–Planckian, negative Wigner functions. In quantum optics [33] negative Wigner functions indicate non–classical features. What are they in our case? When is the spectrum Planckian? What is the temperature? When science answers one question, many more questions arise. We leave it to future readers to find some of the answers.
@@ Acknowledgements
Appendix A Appendix
In this appendix we present an elementary derivation of the relationship between electromagnetic waves and conformally coupled scalar fields [45] propagating in a spatially flat, expanding universe. The line element of the space–time metric, the Robertson–Walker–Friedmann–Lemaître metric, reads [44]:
$\mathrm{d}s^{2}=c^{2}\mathrm{d}t^{2}-a^{2}\mathrm{d}\bm{r}^{2}\,.$ (A.1)
For light rays $\mathrm{d}s=0$, and so we may divide the metric (A.1) by $a^{2}$ with the result
$\mathrm{d}s^{2}_{\mathrm{em}}=\frac{c^{2}}{a^{2}}\,\mathrm{d}t^{2}-\mathrm{d}\bm{r}^{2}=c^{2}\mathrm{d}\tau^{2}-a^{2}\mathrm{d}\bm{r}^{2}$ (A.2)
according to the definition (1) of the conformal time $\tau$. As Maxwell’s equations are conformally invariant [45] the metric (A.2) remains valid for full electromagnetic waves and their fluctuations. Each component $A$ of the vector potential thus satisfies the same wave equation in conformal time $\tau$ and comoving coordinates $\bm{r}$ as in Minkowski space. There we have in Coulomb gauge:
$0=\left(\partial_{\tau}^{2}-c^{2}\nabla^{2}\right)A=\left(a\partial_{t}a\partial_{t}-c^{2}\nabla^{2}\right)A\,.$ (A.3)
Let us define the differential operator
$\Box=\frac{1}{c^{2}}\left(\partial_{t}^{2}+3H\partial_{t}\right)-\frac{\nabla^{2}}{a^{2}}$ (A.4)
in terms of the Hubble parameter (20), and also the quantity
$R=-\frac{6}{c^{2}}\,(\dot{H}+2H^{2})$ (A.5)
and the field
$\phi=\frac{A}{a}\,.$ (A.6)
From definitions (A.4-A.6) follows by elementary calculations
$\bigg{(}\Box-\frac{R}{6}\bigg{)}\phi=\frac{1}{a^{3}c^{2}}\bigg{(}a\partial_{t}a\partial_{t}-c^{2}\nabla^{2}\bigg{)}A=0$ (A.7)
according to the wave equation (A.3). The important point of this relationship is the geometrical meaning of $\Box$ and $R$: $\Box$ is the d’Alembertian [44] $D_{\alpha}D^{\alpha}=\frac{1}{\sqrt{-g}}\partial_{\alpha}\sqrt{-g}g^{\alpha\beta}\partial_{\beta}$ with respect to the Robertson–Walker–Friedmann–Lemaître metric (A.1) with $g=-a^{6}$ and $g^{\alpha\beta}=\mathrm{diag}(1,-a^{-2},-a^{-2},-a^{-2})$. One verifies that $R$ is the corresponding curvature scalar [44]. The resulting wave equation (A.7) is the equation of a conformally coupled scalar field [45]. Q. E. D.